OpenStudy (anonymous):

factorize 12x^2+ x =6

OpenStudy (anonymous):

"factorizing" is not a mathematical property.

OpenStudy (anonymous):

You should use crammer !!

jimthompson5910 (jim_thompson5910):

12x^2+ x = 6 12x^2+ x - 6 = 0 12x^2 + 9x - 8x - 6 = 0 (12x^2 + 9x)+(-8x - 6) = 0 3x(4x + 3)+(-8x - 6) = 0 3x(4x + 3)-2(4x + 3) = 0 (3x-2)(4x + 3) = 0 3x-2 = 0 or 4x+3 = 0 3x = 2 or 4x = -3 x = 2/3 or x = -3/4